# GME: Improving Universal Multimodal Retrieval by Multimodal LLMs

### v1

@@ Abstract
Universal Multimodal Retrieval (UMR) aims to enable search across various modalities using a unified model, where queries and candidates can consist of pure text, images, or a combination of both. Previous work has attempted to adopt multimodal large language models (MLLMs) to realize UMR using only text data. However, our preliminary experiments demonstrate that more diverse multimodal training data can further unlock the potential of MLLMs. Despite its effectiveness, the existing multimodal training data is highly imbalanced in terms of modality, which motivates us to develop a training data synthesis pipeline and construct a large-scale, high-quality fused-modal training dataset. Based on the synthetic training data, we develop the General Multimodal Embedder (GME), an MLLM-based dense retriever designed for UMR. Furthermore, we construct a comprehensive UMR Benchmark (UMRB) to evaluate the effectiveness of our approach. Experimental results show that our method achieves state-of-the-art performance among existing UMR methods. Last, we provide in-depth analyses of model scaling, training strategies, and perform ablation studies on both the model and synthetic data.
@@ 1 Introduction
The growth of multimedia applications necessitates retrieval models that extend beyond traditional text-to-text and text-to-image search (Zhou et al., 2024b). In Universal Multimodal Retrieval (UMR) tasks, both queries and candidates can exist in any modality (Liu et al., 2023c). Compared to addressing this challenge with separate unimodal and cross-modal retrievers in a divide-and-conquer pipeline (Chang et al., 2022), a unified retriever is a more viable option in terms of usability and scalability. Using the dense retrieval paradigm (also known as embedding-based retrieval) (Karpukhin et al., 2020), a unified model can be trained to project inputs from various modalities into a shared embedding space (Zhou et al., 2024b; a; Jiang et al., 2024a). In this space, similarity scores are computed between the embeddings of queries and the retrieval collection, facilitating the efficient ranking of the top- candidates. To achieve this, some previous studies have primarily focused on two approaches: (1) designing feature fusion mechanisms for cross-modal retrievers based on the CLIP architecture (Liu et al., 2023c; Wei et al., 2024), and (2) incorporating visual plugin modules into optimized text embedding models to achieve unified multimodal representations (Zhou et al., 2024b; a).
Recently, researchers have turned to exploring Multimodal Large Language Models (MLLMs) (Liu et al., 2023a; Wang et al., 2024b) in UMR. For example, it is shown that training MLLMs with text data alone can generate universal multimodal embeddings with respectable retrieval performance (Jiang et al., 2024a). However, modality-limited training may fail to fully demonstrate the potential of MLLMs in UMR. We believe that incorporating multimodal data composition (as shown in Figure 1) could further enhance the model performance and generalization. Moreover, visual documents (i.e., document screenshots) are increasingly important in UMR tasks, as they not only simplify the pipelines of diverse Retrieval-Augmented Generation (RAG) applications, but also mitigate information loss during modality conversion (Ma et al., 2024; Faysse et al., 2024b). However, current UMR models primarily target natural images, neglecting support for this scenario (Table 1).
@@
| UniVL-DR (Liu et al., 2023c) | CLIP Feature Fusion | Cross-modal | ✓ | ✗ | ✗ |
| UniIR (Wei et al., 2024) | CLIP Score Fusion | Multimodal | ✓ | ✓ | ✗ |
| BLIP Feature Fusion | |||||
| MARVEL (Zhou et al., 2024b) | Text Encoder+Plugin | Cross-modal | ✓ | ✗ | ✗ |
| VISTA (Zhou et al., 2024a) | Text Encoder+Plugin | Multimodal | ✓ | ✓ | ✗ |
| E5-V (Jiang et al., 2024a) | MLLM | Text-only | ✓ | ✓ | ✗ |
@@
To address the aforementioned challenges, we propose the General Multimodal Embedder (GME), an instruction-based embedding framework utilizing MLLMs as the backbone. GME enables retrieval across various modalities in the unified paradigm, including text, images, visual documents, and fused-modal111We use fuse-modal instead of multimodal to denote the data that contains both text and image to disambiguate from the UMR task. (i.e., image-text composed) contents. Our framework is underpinned by two key techniques: (1) A strategically optimized training data composition for UMR. We categorize UMR tasks into three distinct types: single-modal, cross-modal, and fused-modal (Figure 1). Through extensive experimentation, we analyze how different compositions affect performance (Figure 3) and demonstrate that a balanced mixture of all types yields optimal results. (2) An efficient data synthesis pipeline for large-scale fused-modal training. Recognizing the under-representation of fused-modal data and its potential impact on training effectiveness, we develop a streamlined data synthesis pipeline (§4.2). This approach has successfully generated a comprehensive dataset of 1.1M fused-modal pairs, significantly enhancing our training and model capabilities.
To evaluate the effectiveness of our framework, we compile a comprehensive UMR Benchmark, namely UMRB. This benchmark encompasses tasks from widely recognized retrieval benchmarks in text (Thakur et al., 2021), multimodal (Wei et al., 2024), and visual document retrieval (Faysse et al., 2024b), as well as our newly processed fused-modal retrieval data. We build our models on top of the strong Qwen2-VL series MLLMs (Wang et al., 2024b) and train them on our constructed dataset. Experimental results demonstrate that our model achieves state-of-the-art performance on the UMRB benchmark. Additionally, we perform in-depth analyses on model scaling, training strategies, and ablation of our synthetic data. Our key contributions are:
@@ 2 Related Work
The emergence of Large Language Models (LLMs) has driven significant progress in natural language processing (Brown et al., 2020; OpenAI, 2023b), leading to the development of Multimodal LLMs that extend these capabilities to handle multimodal information. Prominent MLLMs such as GPT-4V (OpenAI, 2023a), LLaVa (Liu et al., 2023a; 2024), Qwen-VL (Wang et al., 2024b), InternVL (Gao et al., 2024; Chen et al., 2024) and MiniCPM-V (Yao et al., 2024) have shown promising advancements in multimodal information understanding and reasoning. Typically, an MLLM consists of an LLM, a vision encoder, and a projector that bridges the two components by transforming raw multimodal inputs into vectors compatible with the LLM (Yin et al., 2023).
@@
Early multimodal retrieval tasks focused on single-modal (Zhao et al., 2024) or cross-modal retrieval (Wang et al., 2016). Recently, the expansion of multimedia applications and multimodal retrieval-augmented generation (RAG) by MLLMs has created a need for unified multimodal retrieval models for complex scenarios. Existing approaches largely utilize pre-trained models such as CLIP (Radford et al., 2021) or BLIP (Li et al., 2022) for multimodal embedding. For instance, UniVL-DR (Liu et al., 2023c) and UniIR (Wei et al., 2024) initially encode images and texts separately using CLIP or BLIP encoders, followed by fusion strategies like score fusion to integrate features from both modalities. Additionally, VISTA (Zhou et al., 2024a) and MARVEL (Zhou et al., 2024b) employ pre-trained text embedding models enhanced with visual plugins to encode composite image-text candidates. However, these methods are typically designed for specific tasks like multimodal document retrieval and lack the flexibility to handle diverse multimodal retrieval tasks.
Concurrent with our work, E5-V (Jiang et al., 2024a) and VLM2VEC (Jiang et al., 2024b) propose fine-tuning MLLMs on single-text (NLI (Gao et al., 2021)) or vision-centric relevance data, demonstrating their transferability to multimodal retrieval. In this paper, we are the first to explore the fine-tuning of an MLLM-based universal multimodal retriever that can address both visual retrieval tasks and maintain strong text-to-text retrieval capabilities. Moreover, we are the first to extend a unified retrieval model to handle not only natural image retrieval but also text-rich image retrieval (Faysse et al., 2024b).
@@
With the advancement of pre-trained Language Models, research in both pure text and Vision-Language Models has focused on building representation models based on these pre-trained language models. In the text retrieval domain, state-of-the-art text embedding models such as Contriver (Izacard et al., 2021), E5 (Wang et al., 2022), GTE (Li et al., 2023), and BGE (Xiao et al., 2024) are all built upon pre-trained language models and have demonstrated impressive generalizability and robust performance in text retrieval tasks. Recently, leveraging LLMs combined with supervised fine-tuning (SFT), researchers have developed unified text representation models that fully utilize the text understanding capabilities of LLMs, resulting in models with enhanced performance and generalization (Wang et al., 2024a; Li et al., 2023; Lee et al., 2024). These models typically process user text inputs through LLMs, using the hidden states from the final transformer layer—either through pooling or by selecting the last token—as the final representation. Inspired by the success of universal text embedding models based on text LLMs, researchers have begun to explore the construction of unified multimodal retrieval models using MLLMs (Jiang et al., 2024a; b). In this paper, we aim to demonstrate through systematic experiments that constructing a truly universal multimodal retrieval model using MLLMs is feasible.
@@ 3 Universal Multimodal Retrieval
Fused-Modal Retrieval: More complicated retrieval tasks involve mixed modalities in queries, candidates, or both. For example, in EVQA (Mensink et al., 2023), both queries and candidates combine text and images.
@@
| Class | Task | Datasets | ||||||||||
TT (16) |
| II (1) | Nights (Fu et al., 2023) | |||||||||||
TI (4) |
@@
TIT (2) | WebQA (Chang et al., 2022) EDIS (Liu et al., 2023b) | ||||||||||
@@
| ITI (2) | FashionIQ (Wu et al., 2021) CIRR (Liu et al., 2021b) | |||||||||||
| ITIT (3) | OVEN (Hu et al., 2023) EVQA (Mensink et al., 2023) INFOSEEK (Chen et al., 2023) |
@@ 3.1 Universal Multimodal Retrieval Benchmark
Based on the aforementioned classification principles, we introduce a new benchmark to comprehensively assess the performance of UMR models. This benchmark comprises 47 evaluation datasets that cover a broad spectrum of multimodal retrieval tasks, and we name it the Universal Multimodal Retrieval Benchmark (UMRB). These evaluation datasets primarily originate from previously constructed datasets tailored for each sub-scenario or sub-task. Specifically, UMRB includes: (1) The BEIR (Thakur et al., 2021) benchmark for text-to-text retrieval scenarios; (2) The M-BEIR (Wei et al., 2024) dataset for vision-centric retrieval scenarios; (3) Additional fused-modal datasets that not cover by M-BEIR; and (4) text-to-visual document search datasets, such as ViDoRe (Faysse et al., 2024b), to extend the coverage of our benchmark and ensure a comprehensive evaluation of model universality. A detailed list of the UMRB datasets is presented in Table 2.
Given the extensive size of UMRB, to expedite our experimental validation and analysis, we have sampled a subset of datasets from each category, constituting a smaller dataset named UMRB-Partial. This subset retains of the total datasets while maintaining evaluation richness. More detailed statistical information about UMRB-Partial can be found in our supplementary materials.
@@ 4 Method
In this section, we present the training framework for developing the General Multimodal Embedder (GME) model using a Multimodal Large Language Model (MLLM). We describe the contrastive learning approach used to train the embedding model. Building on this, we conduct detailed experiments to determine the optimal balance of training data types. Specifically, our experiments demonstrate that diverse data type mixtures significantly enhances the model’s ability to perform retrieval across various modalities. Lastly, recognizing the scarcity of high-quality fused-modal training data, we propose a novel method for automatically synthesizing large-scale, high-quality training data using MLLM.
@@ 4.1 GME: General Multimodal Embedder
We employ a MLLM as the foundation for GME. This model can accept images, text, or image-text pairs as input. Inspired by previous research on text embedding (Li et al., 2023; Wang et al., 2024a), we use the final hidden state of the last token as the representation (or embedding) for the input. Although pre-trained MLLMs possess strong multimodal understanding capabilities, their original training objectives are not optimized for representation learning. Therefore, task-specific fine-tuning (or alignment) is necessary to enhance the model’s representational capacity. Contrastive learning has been shown to effectively train LLMs and MLLMs to produce retrieval embeddings (Li et al., 2023; Jiang et al., 2024a).
@@
The optimization process utilizes the InfoNCE loss function (van den Oord et al., 2018), defined as:
@@
The quality and diversity of negative samples are essential for improving contrastive learning models (Robinson et al., 2020). Inspired by ANCE (Xiong et al., 2021), we employ a two-stage training strategy: (1) Initial Training: We first train the model using randomly selected negative candidates, resulting in Model . (2) Hard Negative Mining and Continue Training: Using , we retrieve the top candidates for each query and select non-relevant candidates from them as hard negatives. We then use these hard negatives to further train , refining it into the final model. This approach ensures the model learns from both easily distinguishable and more challenging examples, thereby enhancing overall performance.
@@
A critical factor in multimodal representation learning is the composition of training data. Although previous studies like (Jiang et al., 2024a) have demonstrated that MLLMs can develop multimodal representation capabilities after being fine-tuned on single-modal data, the effect of data diversity on model performance remains unclear. Therefore, we compare the performance of models trained with different data combinations across various retrieval scenarios within our classification principle. Specifically, we used four types of training data: single-modal (including TT and II), cross-modal (including TVD and TI), fused-modal training data (including ITIT), and a mixed dataset combining the first three types. These different training data types result in a total of six models.
For single-modal data, we utilized the TT dataset from MS MARCO (Nguyen et al., 2016) and the II dataset from ImageNet (Deng et al., 2009), treating images within the same category as positive matches and those from different categories as negatives. For cross-modal data, we employed TI pairs from the LAION (Schuhmann et al., 2022) dataset and TVD pairs from the Docmatix (Laurençon et al., 2024) dataset. For fused-modal data, we use the EVQA (Mensink et al., 2023) dataset (ITIT). For each subcategory, we randomly sampled 100,000 training instances to train the models independently. For the mixed dataset, we uniformly sampled 20,000 instances from each of the five datasets to train the final model, ensuring fair and reliable comparative experimental results. The performance of these six models on the UMRB-Partial test dataset is presented in Figure 3.
The results indicate that: (1) Models trained on single data types excel in corresponding retrieval tasks. For instance, models trained on TT data performed best in text retrieval tasks.333Detail results are shown in the supplementary materials. (2) A balanced mix of different data types enhanced performance across various settings. This suggests that increasing the diversity of training modalities effectively improves the model’s overall retrieval capabilities.
The above analysis highlights the importance of adequately representing each data type in training datasets to develop models that meet the requirements of universal multi-modal retrieval. During data collection, we observed that single-modal and cross-modal data are abundant, with over ten million training instances available. In contrast, fused-modal data remains limited. Common fused-modal training datasets such as EVQA(Mensink et al., 2023), INFOSEEK(Chen et al., 2023), and CIRR (Liu et al., 2021b) collectively contain fewer than one million instances. Additionally, these existing fused-modal datasets cover only a limited range of domains. Thus, efficiently supplementing high-quality fused-modal training data is essential. To address this challenge, we propose leveraging the generative capabilities of LLMs and MLLMs to synthesize additional training data.

### v2

@@ Abstract
Universal Multimodal Retrieval (UMR) aims to enable search across various modalities using a unified model, where queries and candidates can consist of pure text, images, or a combination of both. Previous work has attempted to adopt multimodal large language models (MLLMs) to realize UMR using only text data. However, our preliminary experiments demonstrate that more diverse multimodal training data can further unlock the potential of MLLMs. Despite its effectiveness, the existing multimodal training data is highly imbalanced in terms of modality, which motivates us to develop a training data synthesis pipeline and construct a large-scale, high-quality fused-modal training dataset. Based on the synthetic training data, we develop the General Multimodal Embedder (GME), an MLLM-based dense retriever designed for UMR. Furthermore, we construct a comprehensive UMR Benchmark (UMRB) to evaluate the effectiveness of our approach. Experimental results show that our method achieves state-of-the-art performance among existing UMR methods. Last, we provide in-depth analyses of model scaling and training strategies, and perform ablation studies on both the model and synthetic data.
@@ 1 Introduction
The growth of multimedia applications necessitates retrieval models that extend beyond traditional text-to-text and text-to-image search [75]. In Universal Multimodal Retrieval (UMR) tasks, both queries and candidates can exist in any modality [39]. Compared to addressing this challenge with separate uni-modal and cross-modal retrievers in a divide-and-conquer pipeline [4], a unified retriever is a more viable option in terms of usability and scalability. Using the dense retrieval paradigm (also known as embedding-based retrieval) [25], a unified model can be trained to project inputs from various modalities into a shared embedding space [75, 74, 22]. In this space, similarity scores are computed between the embeddings of queries and the retrieval collection, facilitating the efficient ranking of the top- candidates. To achieve this, some previous studies have primarily focused on two approaches: (1) designing feature fusion mechanisms for cross-modal retrievers based on the CLIP architecture [39, 66], and (2) incorporating visual plugin modules into optimized text embedding models to achieve unified multimodal representations [75, 74].
Recently, researchers have turned to exploring Multimodal Large Language Models (MLLMs) [35, 65] in UMR. For example, it is shown that training MLLMs with text data alone can generate universal multimodal embeddings with respectable retrieval performance [22]. However, modality-limited training may fail to fully demonstrate the potential of MLLMs in UMR. We believe that incorporating multimodal data composition (as shown in Figure 1) could further enhance the model performance and generalization. Moreover, visual documents (i.e. document screenshots) are increasingly important in UMR tasks, as they not only simplify the pipelines of diverse Retrieval-Augmented Generation (RAG) applications, but also mitigate information loss during modality conversion [41, 12]. However, current UMR models primarily target natural images, neglecting support for this scenario (Table 1).
@@
| UniVL-DR [39] | CLIP Feat. Fusion | Cross-modal | ✓ | ✗ | ✗ |
| UniIR [66] | CLIP Score Fusion | Multimodal | ✓ | ✓ | ✗ |
| BLIP Feat. Fusion | |||||
| MARVEL [75] | Text Enc.+Plugin | Cross-modal | ✓ | ✗ | ✗ |
| VISTA [74] | Text Enc.+Plugin | Multimodal | ✓ | ✓ | ✗ |
| E5-V [22] | MLLM | Text-only | ✓ | ✓ | ✗ |
@@
To address the aforementioned challenges, we propose the General Multimodal Embedder (GME), an instruction-based embedding framework utilizing MLLMs as the backbone. GME enables retrieval across various modalities in the unified paradigm, including text, images, visual documents, and fused-modal111 We use fuse-modal instead of multimodal to denote the data that contains both text and image to disambiguate from the UMR task. (i.e. image-text composed) contents. Our framework is underpinned by two key techniques: (1) A strategically optimized training data composition for UMR. We categorize UMR tasks into three types: single-modal, cross-modal, and fused-modal (Figure 1). Through extensive experimentation, we analyze how different compositions affect performance (Figure 3) and demonstrate that a balanced mixture of all types yields optimal results. (2) An efficient fused-modal data synthesis pipeline. Recognizing the under-representation of fused-modal data and its potential impact on training effectiveness, we develop a streamlined data synthesis pipeline (§4.2). This approach has successfully generated a comprehensive dataset of 1.1M fused-modal pairs, significantly enhancing our training and model capabilities.
To evaluate the effectiveness of our framework, we compile a comprehensive UMR Benchmark, namely UMRB. This benchmark encompasses tasks from widely recognized retrieval benchmarks in text [55], multimodal [66], and visual document retrieval [12], as well as our newly processed fused-modal retrieval data. We build our models on top of the strong Qwen2-VL series MLLMs [65] and train them on our constructed dataset. Experimental results demonstrate that our model achieves state-of-the-art performance on UMRB. Additionally, we perform in-depth analyses on model scaling, training strategies, and ablation of our synthetic data. Our key contributions are:
@@ 2 Related Work
The emergence of Large Language Models (LLMs) has driven significant progress in natural language processing [3, 49], leading to the development of Multimodal LLMs that extend these capabilities to handle multimodal information. Prominent MLLMs such as GPT-4V [48], LLaVa [35, 36], Qwen-VL [65], InternVL [7] and MiniCPM-V [71] have shown promising advancements in multimodal information understanding and reasoning. Typically, an MLLM consists of an LLM, a vision encoder, and a projector that bridges the two components by transforming raw multimodal inputs into vectors compatible with the LLM [72].
@@
Early multimodal retrieval tasks focused on single-modal [73] or cross-modal retrieval [61]. Recently, the expansion of multimedia applications and multimodal retrieval-augmented generation (RAG) by MLLMs has created a need for unified multimodal retrieval models for complex scenarios. Existing approaches largely utilize pre-trained models such as CLIP [51] or BLIP [29] for multimodal embedding. For instance, UniVL-DR [39] and UniIR [66] initially encode images and texts separately using CLIP or BLIP encoders, followed by fusion strategies like score fusion to integrate features from both modalities. Additionally, VISTA [74] and MARVEL [75] employ pre-trained text embedding models enhanced with visual plugins to encode composite image-text candidates. However, these methods are typically designed for specific tasks like multimodal document retrieval and lack flexibility to handle diverse multimodal retrieval tasks.
Concurrent with our work, E5-V [22] and VLM2VEC [23] propose fine-tuning MLLMs on single-text (NLI [14]) or vision-centric relevance data, demonstrating their transferability to multimodal retrieval. In this paper, we are the first to explore the fine-tuning of an MLLM-based universal multimodal retriever that can address both visual retrieval tasks and maintain strong text-to-text retrieval capabilities. Moreover, we are the first to extend a unified retrieval model to handle not only natural image retrieval but also text-rich image retrieval [12].
@@
With the advancement of pre-trained Language Models, research in both pure text and Vision-Language Models has focused on building representation models based on these pre-trained language models. In the text retrieval domain, state-of-the-art text embedding models such as Contriver [21], E5 [62], GTE [31], and BGE [68] are all built upon pre-trained language models and have demonstrated impressive generalization and robust performance in text retrieval tasks. Recently, leveraging LLMs combined with supervised fine-tuning (SFT), researchers have developed unified text representation models that fully utilize the text understanding capabilities of LLMs, resulting in models with enhanced performance and generalization [63, 31, 28]. These models typically process user text inputs through LLMs, using the hidden states from the final transformer layer—either through pooling or by selecting the last token—as the final representation. Inspired by the success of universal text embedding models based on text LLMs, researchers have begun to explore the construction of unified multimodal retrieval models using MLLMs [22, 23]. In this paper, we aim to demonstrate through systematic experiments that constructing a truly universal multimodal retrieval model using MLLMs is feasible.
@@ 3 Universal Multimodal Retrieval
Fused-Modal Retrieval: More complicated retrieval tasks involve mixed modalities in queries, candidates, or both. For example, in EVQA [46], both queries and candidates combine text and images.
@@
| Class | Task | Datasets | ||||||
| Single- Modal (17) | TT (16) |
| II (1) | Nights[13] | |||||||
| Cross- Modal (18) | TI (4) |
@@
| Fused- Modal (12) | TIT (2) | WebQA[4] EDIS[37] | ||||||
@@
| ITI (2) | FashionIQ[67] CIRR[38] | |||||||
| ITIT (3) | OVEN[20] EVQA[46] INFOSEEK[6] |
@@ 3.1 Universal Multimodal Retrieval Benchmark
Based on the aforementioned classification principles, we introduce a new benchmark to comprehensively assess the performance of UMR models. This benchmark comprises 47 evaluation datasets that cover a broad spectrum of multimodal retrieval tasks, and we name it the Universal Multimodal Retrieval Benchmark (UMRB). These evaluation datasets primarily originate from previously constructed datasets tailored for each sub-scenario or sub-task. Specifically, UMRB includes: (1) The BEIR [55] benchmark for text-to-text retrieval scenarios; (2) The M-BEIR [66] dataset for vision-centric retrieval scenarios; (3) Additional fused-modal datasets that not cover by M-BEIR; and (4) text-to-visual document search datasets, such as ViDoRe [12], to extend the coverage of our benchmark and ensure a comprehensive evaluation of model universality. A detailed list of the UMRB datasets is presented in Table 2.
Given the extensive size of UMRB, to expedite our experimental validation and analysis, we have sampled a subset of datasets from each category, constituting a smaller dataset named UMRB-Partial. This subset retains of the total datasets while maintaining evaluation richness. More detailed statistical information about UMRB-Partial can be found in Appendix Table 6.
@@ 4 Method
In this section, we present the training framework for developing the General Multimodal Embedder (GME) model. We describe the contrastive learning approach used to train the embedding model. Building on this, we conduct detailed experiments to determine the optimal balance of training data type. Specifically, our experiments demonstrate that diverse data type mixtures significantly enhances the model’s ability to perform retrieval across various modalities. Lastly, recognizing the scarcity of high-quality fused-modal training data, we propose a novel method for automatically synthesizing large-scale, high-quality training data using MLLM.
@@ 4.1 GME: General Multimodal Embedder
We employ a MLLM as the foundation for GME. This model can accept images, text, or image-text pairs as input. Inspired by previous research on text embedding [31, 63], we use the final hidden state of the last token as the representation (or embedding) for the input. Although pre-trained MLLMs possess strong multimodal understanding capabilities, their original training objectives are not optimized for representation learning. Therefore, task-specific fine-tuning (or alignment) is necessary to enhance the model’s representational capacity. Contrastive learning has been shown to effectively train LLMs and MLLMs to produce retrieval embeddings [31, 22].
@@
The optimization process utilizes the InfoNCE loss function [57], defined as:
@@
The quality and diversity of negative samples are essential for improving contrastive learning [53]. Inspired by ANCE [69], we employ a two-stage training strategy: (1) Initial Training: We first train the model using randomly selected negative candidates, resulting in Model . (2) Hard Negative Mining and Continue Training: Using , we retrieve the top candidates for each query and select non-relevant candidates from them as hard negatives. We then use these hard negatives to further train , refining it into the final model. This ensures that the model can learn from both easily distinguishable and more challenging examples, thereby enhancing performance.
@@
A critical factor in multimodal representation learning is the composition of training data. Although previous studies like [22] have demonstrated that MLLMs can develop multimodal representation capabilities after being fine-tuned on single-modal data, the effect of data diversity on model performance remains unclear. Therefore, we compare the performance of models trained with different data combinations across various retrieval scenarios within our classification principle. Specifically, we used four types of training data: single-modal (including TT and II), cross-modal (including TVD and TI), fused-modal training data (including ITIT), and a mixed dataset combining the first three types. These different training data types result in a total of six models.
For single-modal data, we utilized the TT dataset from MSMARCO [47] and the II dataset from ImageNet [10], treating images within the same category as positive matches and those from different categories as negatives. For cross-modal data, we employed TI pairs from the LAION [54] dataset and TVD pairs from the Docmatix [27] dataset. For fused-modal data, we use the EVQA [46] dataset (ITIT). For each subcategory, we randomly sampled 100,000 training instances to train the models independently. For the mixed dataset, we uniformly sampled 20,000 instances from each of the five datasets to train the final model, ensuring fair and reliable comparative experimental results. The performance of these six models on the UMRB-Partial test dataset is presented in Figure 3.
The results indicate that: (1) Models trained on single data types excel in corresponding retrieval tasks. For instance, models trained on TT data performed best in text retrieval tasks.333Detail results are shown in the Appendix Table 8. (2) A balanced mix of different data types enhanced performance across various settings. This suggests that increasing the diversity of training modalities effectively improves the model’s overall retrieval capabilities.
The above analysis highlights the importance of adequately representing each data type in training datasets to develop models that meet the requirements of universal multi-modal retrieval. During data collection, we observed that single-modal and cross-modal data are abundant, with over ten million training instances available. In contrast, fused-modal data remains limited. Common fused-modal training datasets such as EVQA[46], INFOSEEK[6], and CIRR [38] collectively contain fewer than one million instances. Additionally, these existing fused-modal datasets cover only a limited range of domains. Thus, efficiently supplementing high-quality fused-modal training data is essential. To address this challenge, we propose leveraging the generative capabilities of LLMs and MLLMs to synthesize additional training data.